# Twist, splay, and uniform domains in ferroelectric nematic liquid crystals

### v1

@@ VI Conclusions
On the other hand, reorientations of the polarization direction incur elastic energy costs in the form of twists, defect lines, or domain walls. The competition between these two effects creates a variety of domain patterns that depend strongly on the cell thickness and the ionic content. In micron-thin $\mathrm{N}_{\mathrm{F}}$ cells with the same uniform direction of apolar anchoring at the top and bottom plates, the patterns are formed by domains with a uniform polarization that flips by $\pi$ in transition from one domain to the next, see Figs. 1 (a-d), 2(a,b), 4(a), and 7(a). In thicker cells, the neighboring domains, in addition to surface flips, show left- and right-handed $\pi$ twists of polarization around the axis perpendicular to the cell, see Figs. 1(e-j), 3, and 6(a). The lateral extent of these domains in the experiments and in the model $(\lambda_{x}^{*})$ is typically larger than the film thickness and ranges from microns to tens of microns. The length depends on system parameters including elastic constants, polarization density, and the Debye screening length $\kappa^{-1}$. The model derives analytic expressions for the critical concentration of ions above which the domain patterns do not form and for a critical thickness of the cell below which the domains are not twisted. The surface anchoring pattern is another factor that greatly affects the domain structure. In particular, we considered thin cells pre-patterned with a $+1$ radial “aster” defect. Here, the system breaks up into “pie-slice” domains due to the bound charge distribution $\rho=-\nabla\cdot\mathbf{P}\propto 1/r$, decaying with distance $r$ from the defect core. The theoretically predicted number of pie slices $(2n_{\theta}^{*})$ is consistent with the experiment and with the idea that the screening effect in the pre-patterned splay is stronger than in the case of uniform surface anchoring.
@@
In the future, we hope to test the theory more stringently by systematically varying the free ion concentration (and $\kappa$, consequently), which is predicted to have the strongest effect. It would also be interesting to see what happens in a pre-patterned cell with a variety of regions both with vanishing and non-zero $\rho=-\nabla\cdot\mathbf{P}$. We would expect to see “grain boundaries” between different kinds of domain patterning. Finally, an important unexplored question is the nature of the domain wall between $\pi$-twisted domains. The $\pi$-twist soliton considered here (and in Basnet et al. (2022) in more detail) is likely only present for very thin cells which do not have the $\pi$-twist along the $z$-direction. If the domains have counter-rotating twists, as illustrated in Fig. 6(a), then the domains likely consist of other structures, such as pairs of surface disclinations or more discontinuous transitions between polarization orientations Kumari et al. (2023). It would be interesting to investigate if it is also possible to have adjacent domains with the same twist chirality along the $z$ direction. In this case, one would expect a discontinuity in $\mathbf{P}$ orientation along the domain wall and possible uncompensated charge. This would then introduce additional dipolar interactions between pairs of domain walls, leading to an even richer behavior.
Acknowledgements.
We thank J. Selinger and L. Paik for helpful discussions. PK and ODL acknowledge the support of NSF grant DMR-2341830.

### v2

@@ Discussion
In micron-thin $\mathrm{N}_{\mathrm{F}}$ cells with the same uniform direction of apolar anchoring at the top and bottom plates, the patterns are formed by striped domains with a uniform polarization that flips by $\pi$ in transition from one domain to the next, see Figs. 1 a-d, 2a,b, 4a, and 7a. In thicker cells, the stripes show left- and right-handed $\pi$ twists of polarization around the cell normal, see Figs. 1e-j, 3, and 6a. We also considered cells pre-patterned with a $+1$ radial aster defect. Here, the system breaks up into pie-slice domains due to the bound charge distribution $\rho=-\nabla\cdot\mathbf{P}\propto 1/r$, decaying with distance $r$ from the defect core. The theoretically predicted number of pie slices and the stripe width are consistent with the experiment and with the idea that the screening effect in the pre-patterned splay is stronger than in the case of uniform surface anchoring.
@@
In the future, we hope to test the theory more stringently by systematically varying the free ion concentration (and $\kappa$, consequently). It would also be interesting to see what happens in a pre-patterned cell with a variety of regions both with vanishing and non-zero $\rho=-\nabla\cdot\mathbf{P}$. We would expect to see grain boundaries between different kinds of domain patterning. Finally, an important unexplored question is the nature of the domain wall between $\pi$-twisted domains. We found that, given the other assumptions of our model such as a constant $\kappa$, our data were consistent with a domain wall energy density $f_{\pi}\propto h$, corresponding to a fixed width domain wall with elastic deformations spanning the entire cell thickness. It may be possible to have adjacent domains with the same twist chirality along the $z$ direction. In this case, one would expect a discontinuity in $\mathbf{P}$ orientation along the domain wall and possible uncompensated charge, leading to an even richer behavior.
Methods
Materials and sample preparation
We explore an $\mathrm{N}_{\mathrm{F}}$ material abbreviated DIO [2], with a synthesis previously published in [40]. On cooling from the isotropic (I) phase, the phase sequence of DIO is I-$174^{\circ}\mathrm{C}$-$\mathrm{N}$-$82^{\circ}\mathrm{C}$-$\mathrm{SmZ}_{\mathrm{A}}$-$66^{\circ}\mathrm{C}$-$\mathrm{N}_{\mathrm{F}}$-$34^{\circ}\mathrm{C}$-crystal, where N is a regular nematic and $\mathrm{SmZ}_{\mathrm{A}}$ is an antiferroelectric smectic [41]. The sandwich-type cells are bounded by two glass plates with layers of a photosensitive dye, Brilliant Yellow (BY), which shows maximum absorption in the range 400 nm to 550 nm. BY is dissolved in dimethylformamide (DMF) at a concentration 0.5 wt %. The filtered BY-DMF solution is spin-coated onto the substrates at 3000 rpm for 30 seconds and baked for 30 minutes at $90^{\circ}\mathrm{C}$. The spin-coating and baking procedures are performed in a humidity-controlled environment with relative humidity fixed at 0.2.
To achieve apolar planar alignment, the BY-coated assembled cell is exposed to a light beam (light source EXFO X-Cite with a spectral range of 320 to 750 nm) with a linear polarizer for 10 minutes. This irradiation induces bidirectional molecular alignment perpendicular to the polarization axis of the normally incident light. The radial aster pattern of the dye molecules at the substrates is induced by irradiating the substrates through a plasmonic metamask with radial arrangements of nanoslits [11, 12]. The bidirectional apolar anchoring is set by the same light beam that passes through the metamask and acquires local linear polarization orthogonal to the long axis of the nanoslit. The BY molecules realign perpendicularly to the local light polarization, thus the pattern of BY molecules replicates the pattern of nanoslits. To ensure that the surface patterns are the same on the top and bottom plates, these plates are assembled into an empty cell with a preset distance $h$ between them and irradiated by the same light beam. The cell is then filled with DIO in the N phase and then cooled down to the $\mathrm{N}_{\mathrm{F}}$ phase.
Numerical analysis
The estimates of the twist $\tau$ in Fig. 1j and in Fig. 3c are performed by fitting the transmission data to a Jones matrix model using $\tau$ as the fit parameter. The solid lines in these figures represent the fit for the value of $\tau$ given in the legend. The Jones matrix model is implemented in Wolfram Mathematica [42], and the relevant code and description (including all parameters) are available in Ref. [10].
For the stripe width analysis in Fig. 8a, we minimize Eq. (19) with respect to $\lambda_{x}$ using the “FindMinimum” routine in Wolfram Mathematica [42], keeping 500 terms in the summation, which does not appreciably change the result. The expressions for $\lambda_{\mathrm{dw}}$ are varied as described in the Results section in order to get a reasonable match to the data in Fig. 8a.
For the radial sector analysis [curves in Fig. 8(b)], the minimization of Eq. (26) with respect to $n_{\theta}$ is performed numerically, using the Nelder-Mead method implemented in the Scipy Python library [43]. The associated integrals [the one explicitly in Eq. (26) and the one in the definition of $\Xi_{n_{\theta}}(z)$] are evaluated using the method of quadrature, implemented via the Scipy library. The upper bound on the integration in Eq. (26) is taken to be 30 instead of infinite and we checked that lowering this bound did not change the value of the integral within the numerical error of the integrator.
[…]
Acknowledgements
We thank J. V. Selinger, L. Paik, D. Golovaty, and P. Sternberg for helpful discussions. This work was supported by NSF grant DMR-2341830 (O.D.L. and P.K.). After finalizing this work, we learned of a preprint by J. V. Selinger and L. Paik [44] that shows that, when confined to a finite cell, the freely twisting ferroelectric nematic prefers to accommodate an integer number of twist pitches across the cell thickness.